Noel and Joey start from two points which are 590 km apart and travel toward each other. The latter travel at 64 kilometers per hour less than two times the rate of the former. Find their rates if they meet in 5 hours.
Answer by mananth(16946)

let Noel's apeed be x kph
Joey's apeed = 2x-64 kph
...
Since they are travelling towards each other the effective speed = x+2x-64
= 3x-64 kph
...
590/3x-64 = 5
..
590 = 5(3x-64)
590 = 15x-320
15x=910
x= 60.66 Noel's speed
2x-64 = 2*60.66 -64 = 57.32 kph = Joey's apeed
...
m.ananth@hotmail.ca